Circular No. 8212

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ik IN MCG +11-10-56 Further to IAUC 8210, T. Boles reports the discovery of an apparent

supernova (mag 16.6) on unfiltered CCD images taken on Sept. 29.081 and 29.911 UT. The new object is located at α = 8^{h}04^{m}06^{s}.52, δ = +62^{o}59′19′′.8, which is ≈ 9′′.4 west and 18′′.1 north of the center of MCG +11-10-56. SN 2003ik is not present on Boles’ images from Aug. 29 and Sept. 3 (limiting mag 19.5) or on Palomar Sky Survey red or blue plates.

SUPERNOVA 2003il A. Gal-Yam, D. Maoz, E. O. Ofek, D. Poznanski, K. Sharon, E.

Medezinski, and Y. Lipkin, Tel Aviv University; F. Prada, Isaac Newton Group of Telescopes, Instituto de Astrofísica de Canarias; and P. Guhathakurta, University of California at Santa Cruz, report their discovery of an apparent supernova (V ∼ 23) on unfiltered images taken by Ofek on Mar. 26 with the 2.5-m Nordic Optical Telescope (+ ALFOSC imaging spectrograph) at La Palma. The new object is located at α = 16^{h}35^{m}47^{s}.5, δ = +66^{o}13′26′′.9 (equinox 2000.0), in the field of the galaxy cluster Abell 2218 (z = 0.175). SN 2003il was invisible in images obtained in Apr. and June 2002 (limiting mag V ∼ 25.0). P. Challis, T. Matheson, and R. Kirshner report that this object is confirmed in V, R, and I frames obtained with the 6.5-m MMT telescope (+ MINICAM) on 2003 May 5. Gal-Yam and Sharon add that inspection of deep archival images of Abell 2218, obtained with the Hubble Space Telescope (broad-V F606W filter in Jan. 2000; broad-R F702W filter in Sept. 1994), reveals no sources near the location of SN 2003il (limiting magnitude equivalent to V = 27). These strong limits on the luminosity of possible host galaxies suggests that this supernova may be yet another example of the class of intergalactic supernovae in clusters (Gal-Yam et al. 2003, A.J. 125, 1087). The distance of SN 2003il from the cluster core is 76′′ (233000 pc, assuming a flat cosmology and a Hubble constant of 70 km/s/Mpc). Using the colors measured from the imaging data obtained by Challis et al. (I = 23.4 ± 0.12, R = 23.6 ± 0.16, V = 24.5 ± 0.16), combined with the color-classification tools of Poznanski et al. (2002, PASP 114, 833), Gal-Yam et al. find that SN 2003il is consistent with a type-Ia supernova at the redshift of the cluster, discovered some months after peak magnitude, which is also possibly somewhat underluminous for its age (by ∼ 1 mag). They add that the supernova colors can also be explained by a higher-redshift event (i.e., behind the cluster Abell 2218) exploding in an undetected low-luminosity dwarf host.

2003 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green